OpenStudy (anonymous):

trying to solve for x ....x-10 squareroot ofx +9=0

OpenStudy (anonymous):

x-10\[\sqrt{x}+9\]

OpenStudy (anonymous):

anybody?

OpenStudy (anonymous):

not clear \[x-10\sqrt{x+9}=0\]?

OpenStudy (anonymous):

\[x-10\sqrt{x}+9=0\]

OpenStudy (anonymous):

idk why it comes out like that but its x-10 and then suare root of x and then +9 ?